Complete integral of pq=xy


1
Expert's answer
2021-01-03T14:43:42-0500

Given ,

"\;\;\;\;pq=xy\\
\Rightarrow \frac{p}{x}=\frac{y}{q}"

Let

"\;\;\;\;\frac{p}{x}=\frac{y}{q}=k\\
\Rightarrow p=xk\;\;\;and, \;\;\;q=\frac{y}{k}"


Now,

"dz=pdx+qdy=xkdx+\frac{y}{k}dy"

Integrating both sides,

"\;\;\;\int dz=k\int xdx+\frac{1}{k}\int ydy\\\;\\
\Rightarrow z=\frac{kx^2}{2}+\frac{y^2}{2k}+c' \\
[c'\;integrating\;constant]\\\;\\
\Rightarrow2kz=k^2x^2+y^2+c\\
[c=2c'k]"


"\therefore" So, the required equation :-

"\boxed{2kz=k^2x^2+y^2+c}"